OpenStudy (anonymous):

What are the solutions of the system? y = x^2 + 4x - 5 y = 2x - 2 A. (–3, –8) and (1, 0) B. (–3, –8) and (–1, –4) C. (–3, 0) and (1, –8) D. no solution A, B, C, or D? Please explain :)

OpenStudy (anonymous):

Well, they are both equal to y so therefore they are both equal to each other. What do you get when you combine them and then factor?

OpenStudy (anonymous):

I don't get anything, I don't know how to do it.

OpenStudy (anonymous):

Equate both the things above..

OpenStudy (anonymous):

See: If : a = 3y + 7 a = 4y + 2 Then : to equate them means " 3y + 7 = 4y + 2..

OpenStudy (anonymous):

x^2+4x-5=2x-2

OpenStudy (anonymous):

Yes @malical is right. Now you will get a quadratic equation, by solving it you will get two values of x..

OpenStudy (anonymous):

\[x^2+4x-5=2x-2\] \[x^2 + 2x - 3 = 0\]

OpenStudy (anonymous):

Can you factorize it @zigzagoon2000 ??

OpenStudy (anonymous):

You should try or participate @zigzagoon2000 After all it is your question..

OpenStudy (anonymous):

I apologize, I was in the other room. But I do not how to do any of that.

OpenStudy (anonymous):

You don't know to get to that point or you don't know how to factor?

OpenStudy (anonymous):

Both

OpenStudy (anonymous):

Okay. Can we accept that y=y and therefore x^2+4x-5=2x-2?

OpenStudy (anonymous):

I guess so, I mean like I said: I don't know how to do any of this stuff.

OpenStudy (anonymous):

They gave you two equations y=x^2+4x-5 and y=2x-2. These mean that when you put in say x =0 you get y=-5 and y=-2. Or the points: (0,-5) and (0,-2). What this problem is asking for is for you to find a value x such that the y's are equal. So what we want to do is to solve for x. So: y=x^2+4x-5 y=2x-2 In this system, both of these statements are true. And because we want the y's to be equal, we can say that y=y. Therefore, the equations with the x's attached to the are also equal. y=y=x^2+4x-5=2x-2 The important part of this being that: x^2+4x-5=2x-2 Our next step to solve this is focused on factoring these. To do that we first need to get everything to to left side so that the right side is 0. x^2+4x-5-2x+2=0 Then simplify that: x^2+2x-3=0 Are you okay up to this point? We can relook at any part.